Estimate the number of gallons in Earth's oceans (in units of 1e21 gallons).

If one's estimate was 1e22 gallons, one would enter 10 as 1e22 / 1e21 = 10.

If one's estimate was 1e20 gallons, one would enter 0.1 as 1e20 / 1e21 = 0.1.
Answer #1

Earth's Ocean water volume is 3.52×10^20 gallons

As unit is 10^21, so ocean volume = (3.52×10^20)/10^21= 0.352